OpenStudy (anonymous):

What is the volume of a prism where the base is a 45-45-90 triangle with a leg of 5 in. when the height is 1.8 in.? a.19.32 in3 b.21.99 in3 c.22.5 in3 d.18.

OpenStudy (anonymous):

D is supposed to be 18.77

OpenStudy (raden):

the area of base = 1/2 * 5 * 5 = 25/2 = 12.5

OpenStudy (anonymous):

But that isnt an answer

OpenStudy (raden):

so, volume that prism = area base * its height = 12.5 * 1.8 = 22.5

OpenStudy (anonymous):

Ohh okay thank you

OpenStudy (raden):

very welcome ;)